A man driving a truck at  a constant speed of 20 m/s suddeny saw a parked car ahead him by 95m. He could apply brakes after some time to produce a retardation of 2.5m/s^2 .If an accidentwas just avoided find his reaction time 

Grade:11

1 Answers

Jobin C J
6 years ago
s=95-20∆t,v²=u²-2as,400=2×2.5×(95-20∆t),400/2×2.5 = 95-20∆t,80 = 95-20∆t,∴20∆t=15∆t = 3/4 =0.75 SecondsJobin C JXII